OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=6tan(x) at point (pi/4, 6)

OpenStudy (e.mccormick):

Using the limit definition or do you know derivatives?

OpenStudy (e.mccormick):

I ask, because when people get tangent line questions, they may still be in the limit version. But you are using a transcendental equation so... perhaps, perhaps not. \[f'(x)=\lim_{\Delta x\to 0}\frac{f(x+\Delta x)-f(x)}{\Delta x}.\]

OpenStudy (anonymous):

here dy/dx = 6sec^2(x) put x=pi/4 u'll get slope as m=12 and u have a point (pi/4,6) so u have slope and a point i think u can get equation of tangent

OpenStudy (deogwa88):

use the derivative of the original equation, plug in the x value to get M Use the form (y-y1)=M(x-x1) y1 is the y cordinate i.e. 6 and x1 is the x cordinate pi/4